Pete and​ Jessica, on the advice of their​ next-door neighbor, recently purchased 600 shares of a​ small-capitalization Internet​ stock, trading at $ 78.91 per share. Their neighbor told them that the stock was a​ "real money​ maker" because it recently had a​ two-for-one stock split and would probably split again soon. Even​ better, according to the​ neighbor, the company was expected to earn $ 1.49 per share and pay a $0.27 dividend next year. Pete and Jessica have so far been less than impressed with the​ stock's performancelong dash —the stock has underperformed the​ S&P 500 Index this year. Pete and Jessica have come to you for some independent advice. Required: a. Assuming that the stock actually splits two for​ one, how many shares will Pete and Jessica​ own? What will be the market value of their stock after the​ split? How will the split affect the value of their​ holdings? Was their neighbor correct in thinking that the stock split made the stock a​ "real money​ maker"? b. Using the information​ provided, calculate the​ stock's P/E ratio. Would you classify this investment as a growth or value​ stock? c. Since​ Pete, in​ particular, is worried about the price of the​ stock, explain to him how and why corporate earnings are so important in the valuation of common stocks.
Answer #1

a]

Shares owned after split = Shares owned before split * split ratio

Shares owned after split = 600 * 2 = 1,200

share price after split = share price before split / split ratio

share price after split = $78.91 / 2 = $39.455

market value of their stock after the​ split = share price after split * Shares owned after split

market value of their stock after the​ split = $39.455 * 1,200 = $47,346

market value of their stock before the​ split = share price before split * Shares owned before split

market value of their stock before the​ split = $78.91 * 600 =  $47,346

The value of their holdings are not affected by the split.

No, their neighbor was not correct in thinking that the stock split made the stock a​ "real money​ maker". This is because the value of their holdings are not affected by the split. A stock is split to make it more liquid in the market. It does not make the stock a money maker

b]

P/E ratio = stock price / EPS

P/E ratio = $78.91 / $1.49 = 52.96

Growth stocks usually have high P/E ratio. The P/E ratio of this stock is high. It can be classified as a growth stock

c]

Corporate earnings are important in the valuation of common stocks because :

  • Earnings represent return on the capital invested. Any investor invests with the expectation of earning a return on their investment.
  • Earnings generate cash flows which are available for distribution to shareholders as dividends